OpenStudy (anonymous):

find four consecutive even integers such that if the sum of the first and third is multiplied by 2, the result is 20 less than two times fourth.

OpenStudy (experimentx):

2x, 2(x+1), 2(x+2), 2(x+3) 2*(2x+2(x+2)) + 20 = 2*2(x+3) find the value of x

OpenStudy (anonymous):

thank you

OpenStudy (experimentx):

yw

OpenStudy (campbell_st):

the condition of the 1st and 3rd number being a multiple of 2 doesn't need to be considered as the sum of any 2 even numbers is even, and divisible by 2 let the even numbers be x , x + 2, x + 4 and x + 6 then setting up the equation x + x+2 + x + 4 + x + 6 = 2(x +6) - 20 gives 4x + 12 = 2x - 8 x = -10 the numbers are -10, -8, -6, -4